CAT Chat Minutes

July 12, 2004

 

Information and Follow-up
There was no dump as a result of the earthquake that took place in Ottawa, Illinois, however we have seen that there was a dump when the earthquake occurred in Alaska recently. Why didn't the recent local earthquake trip the beam?

This is interesting; we could check the record log to see just exactly the effect of each earthquake was to our location here at the APS.  G. Decker may have those numbers and we can take a look at them and report findings.

General Information
Machine Stats
 
Delivered beam:   972 hrs.
   Percentage of scheduled time:   98.46
   Downtime during period:   15.19 hrs.
   Mean time between faults:   51.18 hrs.

Faults:

# 26 07/05 08:00 to 07/07 19:44 59.74

Intentional dump

0.36 corrupted 0+324 fill pattern, dumped, refill

#29 07/07 20:06 to 07/11 10:25 86.24

Power sag / ComEd

0.59 power sag to 84%, restored systems, refilled SR

*Interruption of delivered beam between 07/09 17:29 to 07/09 17:34 users shutters closed due to an ACIS inhibit

Fill ongoing

0.00  

#30 07/11 10:55 to 07/12 09:05 22.16

Fill ongoing

-0.00

Beamline Downtime Summary by APS Groups, 06/29/2004 to 07/12/2004:

DIN

Date

Beamline

Status

Duration

Group

Description

Resolution

424

07/01/04

09-ID

Open

2 hr 49 min

ASD-VAC

Cooling water for a beamline mask in the FOE (downstream of the diff. pump and upstream of the BIV) was turned off during the shutdown for the replacement of the BIV   (Work Request #15116). It was noticed that the mask has been running without cooling water. The mask was not part of the BL-EPS . .

425

07/02/04

18-ID

Closed

22 min

ASD-VAC

PSS Fault #67 PS2_NO_SWITCH occurred.

ASD-VAC will affect repairs on PS2 during the next shutdown.
J. Gagliano will monitor the status of this shutter for the duration of the run.

427

07/11/04

08-ID

Open

16 min

Fault 208 9DMS1_NOSWITCH. Shutter was open at the time of the fault, but there were no experimenters present. Reset per standing order.

428

07/11/04

08-ID

Open

50 min

Fault 67 PS2_NO_SWITCH-S. Contacted PSS ( Friedman ) and reset fault as per standing order.

Although not confirmed, DIN 427 of 8ID is due to a cylinder problem. A timer relay in 18ID will be switched out which very well may take care of the 8ID problem.

LEUTL:  Roger projected a viewgraph which displayed an elevation view of the LEUTL radiation stop.  This stop allows acess to the LEUTL shielded enclosure when the upstream "pitching" magnets are turned off so that the booster electron beam is directed into the x-ray ring or into the booster BTS beam dump.  The cylinders on the radiation stop were replaced during the last outage.  It was discovered, last Friday, that the cylinders and switches were installed improperly.  When the stop indicated "open" it was in a "closed" state, and vice-versa.  There were three levels of protection provided by the two pitching magnets and the rad monitor in the LEUTL enclosure that still would prevent personnel accessing the LEUTL enclosure from being exposed to radiation.  This incident is presently being investigated by the APS.

Safety E. Chang: The construction fence has been erected and the gates will remain open until Bluff Road (by New Brunswick lab) is finished being resurfaced.  

Bruce Glagola, AOD Safety, will be on vacation and unreachable from July 17-25 and and August 1-8, 2004.

Mail Lists:   If you are interested in receiving automatic e-mails regarding APS beam status, remember to subscribe to the CAT INFO mail list.

Other:  In order to reduce background contamination on the CU XBPMs , Glenn Decker and the OAG group have finalized a Decker distortion configuration that will be put into place on all CU lines in the future.